Can you find the value of n and c to make the systems of equations true?
c(n)=-3n-6
c(n)=n-4

Answer by TimothyLamb(4379)

c(n) = -3n - 6
c(n) = n - 4
---
-3n - 6 = n - 4
-4n = 2
---
n = -1/2 = -0.5
c(n) = -9/2 = -4.5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php